### v1

# Electron charge coherence on a solid neon surface

Abstract
Recent experiments show $\sim$0.1 ms coherence time for a single electron charge qubit on a solid neon surface. This remarkably long coherence time is believed to result from the intrinsic purity of solid neon as a qubit host. In this paper, we present theoretical studies on the decoherence mechanisms of an electron’s charge (lateral motional) states on solid neon. At the typical experimental temperature of $\sim$10 mK, the two main decoherence mechanisms are the phonon-induced displacement of neon surface and phonon-induced modulation of neon permittivity (dielectric constant). With a qubit frequency increasing from 1 GHz to 10 GHz, the charge coherence time decreases from about 366 s to 7 ms and from about 27 s to 0.3 ms, respectively, limited by the two mechanisms above. The calculated coherence times are at least one order longer than the observed ones at $\sim$6.4 GHz qubit frequency, suggesting plenty of room for experimental improvement.
I Introduction
Among solid-state qubits [1, 2], electron charge qubits offer promising prospects to serve as building blocks in quantum computing, due to their inherent simplicity in design [3, 4], fabrication [5], control [6, 7, 8], and readout [9, 4, 10, 11, 12]. However, a notable drawback of electron charge qubits is their generally short coherence times, attributed to their sensitivity to environmental charge noises. The pursuit of novel electron charge qubits with extended coherence times is highly attractive in quantum information science and technology.
Helium (He) and neon (Ne), noble-gas elements with extreme chemical inertia, spontaneously condense into ultrapure quantum liquids and solids at low temperature and pressure [13]. Compared with classical liquids and solids, condensed He and Ne have negligible charged excitations or spinful isotopes (3He and 21Ne). They offer a vacuum-like environment with minimal charge and spin noises for electron qubits [14, 13]. When an excess electron approaches a liquid He or solid Ne from a vacuum, it can form surface states by two effects. First, there is a repulsive barrier of approximately 1 eV resulting from the Pauli exclusion between the excess electron and the shell electrons in He and Ne atoms. Second, there is an attractive well from the polarization-induced image charge inside the liquid or solid [15, 16, 17]. A surface electron can be further trapped in a quantum-dot structure by on-chip electrodes and coupled with microwave photons in a superconducting quantum circuit [18, 19, 20].
In the past few years, we successfully realized the electron-on-solid-neon (eNe) motional (charge) qubits [21] and obtained charge coherence times up to 0.1 ms [22]. Despite already exceeding the coherence times of traditional superconducting and semiconductor charge qubits by several orders [9, 23], the current eNe charge coherence times are probably limited by the imperfect surface and redundant electrons trapped on it [24, 25]. With optimized growth of solid Ne and refined electron loading procedure, the eNe charge coherence times may be significantly improved.
In this paper, we present a theoretical investigation of the intrinsic decoherence mechanisms of the motional (charge) states of a single electron on a perfect solid Ne surface. We calculate the corresponding relaxation and coherence times, which provide the upper limit of the eNe charge coherence times in this system. The system is illustrated in Fig. I, which mimics the thin-film structures of typical fabricated devices. A single electron hovers above the surface of a solid Ne film with a thickness denoted by $w$. The devices are typically superconducting circuits made of aluminum (Al) or niobium (Nb) film with a thickness denoted by $t$, deposited on low-microwave-loss substrates such as intrinsic silicon (Si) or sapphire (Al2O3). The solid Ne layer can be grown on top of both the metal electrode and the bare substrate.
Fig. 1. : Schematic illustration of the electron-on-solid-neon (eNe) system, featuring an electron floating in the vacuum above a solid Ne film. The solid neon film can cover both the metal electrodes and the dielectric substrate of the quantum circuit device. There are longitudinal (acoustic) phonons in the bulk and their induced displacement on the surface. The electron’s wavefunction, represented by a red cloud, resides at a mean distance of $d\approx 2$ nm from the surface and extends along the surface over a scale of 100 nm. The diagram shows a harmonic ground-state wavefunction (pink curve) with a Gaussian width $a_{x}$.
II Surface electronic states
In this paper, we adopt the simplest model to describe the surface electronic states [26]. We approximate the Pauli barrier to be infinitely high, which forces the electron wavefunction to vanish inside solid Ne, $z\leq 0$. We also assume the solid Ne to be thick, $w\gtrsim 100$ nm, so the influence from the metal electrode and substrate on the wavefunction is negligible. The image potential in the vertical ($z$) direction is
$V_{\text{img}}(z)=-\frac{\mathit{\Lambda}}{z},\quad(z>0),$ (1)
where $\mathit{\Lambda}$ is the effective Coulomb interaction parameter,
$\mathit{\Lambda}=\frac{e^{2}}{4}\frac{\epsilon_{\scriptscriptstyle\text{Ne}}-1}{\epsilon_{\scriptscriptstyle\text{Ne}}+1},$ (2)
with $e$ being the elementary charge and $\epsilon_{\scriptscriptstyle\text{Ne}}=1.244$ being the dielectric constant of uniform solid neon.
The electron’s eigenenergies are hydrogenic,
$E_{n}=-\frac{R}{n^{2}},\quad n=1,2,3,\dots$ (3)
with the effective Rydberg energy,
$R=\frac{\hbar^{2}}{2m_{e}r_{\scriptscriptstyle\text{B}}^{2}}=1.611\times 10^{-14}\,\text{erg}=116.7\,\text{K}.$ (4)
in which $m_{e}$ is the electron mass and $r_{\scriptscriptstyle\text{B}}$ is the effective Bohr radius,
$r_{\scriptscriptstyle\text{B}}=\frac{\hbar^{2}}{\mathit{\Lambda}m_{e}}=1.947\,\text{nm},$ (5)
They are related by $\mathit{\Lambda}=2Rr_{\scriptscriptstyle\text{B}}$. The ground-state wavefunction is
$\psi_{1}(z)=\frac{2}{\sqrt{r_{\scriptscriptstyle\text{B}}}}\frac{z}{r_{\scriptscriptstyle\text{B}}}{\rm e}^{-z/r_{\scriptscriptstyle\text{B}}}.$ (6)
At the typical experimental temperature, $T=10$ mK, the electron’s vertical motion perpendicular to the Ne surface is effectively frozen in the ground state $\left|1_{z}\right\rangle$, whereas its lateral motion parallel to the $xy$ plane is confined by on-chip electrodes. We approximate the trapping potential in each of the lateral dimensions to be harmonic, with level spacing $\hbar\omega_{x,y}$. Thus, the Hamiltonian ^$\hat{\mathcal{H}}_{\text{el}}$ for a single electron is given by
^$\hat{\mathcal{H}}_{\text{el}}=-\frac{\hbar^{2}}{2m_{e}}\nabla^{2}+V_{x}(x)+V_{y}(y)+V_{z}(z),$ (7)
with
$V_{z}(z)=V_{\text{img}}(z)+e\mathcal{E}_{z}z,\quad z>0,$ (8)
where $\mathcal{E}_{z}$ is an electric field that presses the electron against the Ne surface, and
$V_{x}(x)=\frac{1}{2}m_{e}\omega_{x}^{2}x^{2},\quad V_{y}(y)=\frac{1}{2}m_{e}\omega_{y}^{2}y^{2},$ (9)
are lateral harmonic potentials in $x$ and $y$, respectively.
The in-plane motional states of the electron are denoted as $\left|n_{x},n_{y}\right\rangle$ with $n_{x},n_{y}=0,1,2,\dots$ being the quantum numbers of quantum harmonic oscillators in $x$ and $y$, respectively. Although in this model an electron’s wavefunction entirely resides in the vacuum, phonon excitations in solid Ne can still decohere electron’s emotional states, as discussed below.
III Decoherence via photons
Before talking about phonons, let us first talk about photons which set the upper bound of the relaxation and coherence times in this system.
III.1 Spontaneous emission in a vacuum
The electric dipole transition can radiate photons into the surrounding vacuum. The spontaneous emission rate can be described as [27]
$\mathit{\Gamma}^{\left(\text{vac}\right)}=\frac{\left|d_{x}\right|^{2}}{3\hbar c^{3}}\omega_{x}^{3}\equiv\frac{1}{T_{1}^{\left(\text{vac}\right)}}.$ (10)
Here, $d_{x}=\sqrt{2}ea_{x}$ is the electric dipole moment in $x$ with the confining length $a_{x}=\left(\hbar/m\omega_{x}\right)^{1/2}$, and $f_{x}=\omega_{x}/2\pi$ is the transition frequency. The rate of radiative decay, as given by Eq. (10), determines the relaxation time $T_{1}^{\left(\text{vac}\right)}$. For our typical choice of $f_{x}=6.4$ GHz, the corresponding $T_{1}^{\left(\text{vac}\right)}$ is 999999 s. This gives a remarkably long upper bound of the coherence time $T_{2}^{\left(\text{vac}\right)}$ , ref [28],
$T_{2}^{\left(\text{vac}\right)}\leq 2T_{1}^{\left(\text{vac}\right)}=198\,\text{s}.$ (11)
III.2 Spontaneous emission in a superconducting cavity
In practice, the electron is often situated in an on-chip superconducting resonator. The resonator alters the spontaneous emission rate due to the modified electromagnetic environment, known as the Purcell effect[29]. For a simple two-level electric dipole coupled to a single-mode resonator with the resonator frequency $f_{\text{r}}=\omega_{\text{r}}/2\pi$ through nonresonant coupling (i.e., $g\ll\left|\omega_{x}-\omega_{\text{r}}\right|$), the modified spontaneous emission rate can be described as [22]
$\mathit{\Gamma}^{\left(\text{cav}\right)}=\frac{g^{2}\kappa}{\left(\omega_{x}-\omega_{\text{r}}\right)^{2}}\text{,}$ (12)
Here, $\kappa$ is the resonator linewidth (leakage rate). For example, with $g/2\pi\sim 5$ MHz and $\kappa/2\pi\sim 0.5$ MHz [22], if $\left|f_{x}-f_{\text{r}}\right|\sim 500$ MHz, the relaxation time of the electron in the cavity is shortened to $T_{1}^{\left(\text{cav}\right)}=32$ ms. As a result, the coherence time is limited with an upper bound of $T_{2}^{\left(\text{cav}\right)}\leq 64$ ms, which is much shorter than the upper bound in vacuum but still longer than the coherence times limited by other mechanisms below. It is worthwhile to note that the Purcell limited radiative decay can be suppressed by introducing Purcell filter [30] or other readout techniques.
IV Decoherence via phonons
The electron’s lateral motion outside the solid Ne can couple with the longitudinal (acoustic) phonons inside solid Ne. Consequently, the electron’s charge states can undergo relaxation and dephasing via emitting phonons.
The displacement field $\bm{u}(\bm{r},t)$ of solid Ne due to the longitudinal (acoustic) phonons, in a large space of volume $V$ and undisturbed mass density $\rho_{\scriptscriptstyle\text{Ne}}=1.444$ g cm-3, can be written as
$\bm{u}(\bm{r},t)=\sum_{\bm{q}}\frac{\bm{q}}{q}\sqrt{\frac{\hbar}{2V\rho_{\scriptscriptstyle\text{Ne}}\omega_{q}}}(a_{\bm{q}}{\rm e}^{-{\rm i}\omega_{q}t}-a^{\dagger}_{-\bm{q}}{\rm e}^{+{\rm i}\omega_{q}t}){\rm e}^{{\rm i}\bm{q}\cdot\bm{r}},$ (13)
where $\bm{r}=(x,y,z)$ and $\bm{q}=(q_{x},q_{y},q_{z})$ are the 3D positions and wavevectors, respectively; $a_{\bm{q}}$ and $a^{\dagger}_{\bm{q}}$ and are the phonon annihilation and creation operators of the plane-wave mode with the wavevector $\bm{q}$; $\omega_{q}=c_{\scriptscriptstyle\text{Ne}}q$ is the dispersion relation that only depends on the magnitude (not direction) of the wavevector, $q=|\bm{q}|=\sqrt{q_{x}^{2}+q_{y}^{2}+q_{z}^{2}}$, $c_{\scriptscriptstyle\text{Ne}}=1.133\times 10^{5}$ cm s-1 is the longitudinal sound velocity; $\bm{q}/q$ is the unit vector in the longitudinal wavefront direction.
The velocity field $\bm{v}(\bm{r},t)$ can be derived as
$\bm{v}(\bm{r},t)=\sum_{\bm{q}}\frac{\bm{q}}{{\rm i}q}\sqrt{\frac{\hbar\omega_{q}}{2V\rho_{\scriptscriptstyle\text{Ne}}}}(a_{\bm{q}}{\rm e}^{-{\rm i}\omega_{q}t}+a^{\dagger}_{-\bm{q}}{\rm e}^{+{\rm i}\omega_{q}t}){\rm e}^{{\rm i}\bm{q}\cdot\bm{r}}.$ (14)
The acoustic velocity potential $\mathit{\Phi}(\bm{r},t)$ is
$\mathit{\Phi}(\bm{r},t)=\sum_{\bm{q}}\frac{1}{q}\sqrt{\frac{\hbar\omega_{q}}{2V\rho_{\scriptscriptstyle\text{Ne}}}}(a_{\bm{q}}{\rm e}^{-{\rm i}\omega_{q}t}+a^{\dagger}_{-\bm{q}}{\rm e}^{+{\rm i}\omega_{q}t}){\rm e}^{{\rm i}\bm{q}\cdot\bm{r}},$ (15)
which satisfies
$\bm{v}=-\nabla\mathit{\Phi}.$ (16)
The density variation field $\varrho(\bm{r},t)\equiv\rho(\bm{r},t)-\rho_{\scriptscriptstyle\text{Ne}}$, where $\rho(\bm{r},t)$ is the actual density field, is
$\varrho(\bm{r},t)=\sum_{\bm{q}}\frac{1}{{\rm i}}\sqrt{\frac{\hbar\omega_{q}\rho_{\scriptscriptstyle\text{Ne}}}{2Vc_{\scriptscriptstyle\text{Ne}}^{2}}}(a_{\bm{q}}{\rm e}^{-{\rm i}\omega_{q}t}-a^{\dagger}_{-\bm{q}}{\rm e}^{+{\rm i}\omega_{q}t}){\rm e}^{{\rm i}\bm{q}\cdot\bm{r}},$ (17)
which satisfies the acoustic continuity equation,
$\partial_{t}\varrho(\bm{r},t)+\rho_{\scriptscriptstyle\text{Ne}}\nabla\cdot\bm{v}(\bm{r},t)=0,$ (18)
[…]
$\varrho(\bm{r}_{\scriptscriptstyle\parallel},z)$ $=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\frac{1}{{\rm i}}\sqrt{\frac{\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}\rho_{\scriptscriptstyle\text{Ne}}}{Vc_{\scriptscriptstyle\text{Ne}}^{2}}}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}}\sin(q_{z}z)$
×$\times(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}),$ (19)
the form of which ensures that the density variation vanishes at the Ne-vacuum interface, $z=0$, where pressure is zero. The summation in Eq. (19) only includes $q_{z}>0$ to ensure that the modes are counted correctly. Likewise, the acoustic velocity potential is
$\mathit{\Phi}(\bm{r}_{\scriptscriptstyle\parallel},z)$ $=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\sqrt{\frac{\hbar c_{\scriptscriptstyle\text{Ne}}}{V\rho_{\scriptscriptstyle\text{Ne}}q}}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}}\sin(q_{z}z)$
×$\times(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}+\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}).$ (20)
And the displacement field is
$\bm{u}(\bm{r}_{\scriptscriptstyle\parallel},z)$ $=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\frac{\bm{q}}{q}\sqrt{\frac{\hbar}{V\rho_{\scriptscriptstyle\text{Ne}}\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}}}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}}\cos(q_{z}z)$
×$\times(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}).$ (21)
There are two major electron-phonon coupling mechanisms in this system [31, 32]: 1. Phonon-induced displacement of the neon surface, which interacts with the electron via the Pauli repulsion and the varied image potential. 2. Phonon-induced modulation of the neon density and hence the dielectric constant, which interacts with the electron via the varied image potential.
The total Hamiltonian consists of the electron part ^$\hat{\mathcal{H}}_{\text{el}}$ as given above, the phonon part ^$\hat{\mathcal{H}}_{\text{ph}}$, and the electron-phonon coupling part ^$\hat{\mathcal{H}}_{\text{el-ph}}$,
^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\text{el}}+\hat{\mathcal{H}}_{\text{ph}}+\hat{\mathcal{H}}_{\text{el-ph}},$ (22)
where
^$\hat{\mathcal{H}}_{\text{ph}}=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}a^{\dagger}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}a_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}},$ (23)
and
^$\hat{\mathcal{H}}_{\text{el-ph}}=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z){\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}}(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}).$ (24)
The Hamiltonian is written in the single electron position representation and the second-quantized phonon fields on the plane-wave bases and ignores the zero-point energy. The operators $\hat{a}_{\bm{q}}=\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$ and $\hat{a}_{\bm{q}}^{\dagger}=\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}^{\dagger}$ are the same phonon annihilation and creation operators as before except being written explicitly in terms of $\bm{q}_{\scriptscriptstyle\parallel}$ and $q_{z}$. $\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)$ is the coupling strength between an electron at the position $(\bm{r}_{\scriptscriptstyle\parallel},z)$ and a plane-wave phonon of the wavevector $(\bm{q}_{\scriptscriptstyle\parallel},q_{z})$. $\omega_{\bm{q}}=\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$ is the same phonon dispersion relation as before, $\omega_{\bm{q}}=\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}=c_{\scriptscriptstyle\text{Ne}}q=c_{\scriptscriptstyle\text{Ne}}\sqrt{q_{\scriptscriptstyle\parallel}^{2}+q_{z}^{2}}$.
The relaxation rate corresponding to electron scattered by a single phonon can be calculated from the Fermi golden rule,
$\mathit{\Gamma}$ $=\frac{2\pi}{\hbar}\left|\langle n_{\text{ph}}^{f}|\langle 0|\hat{\mathcal{H}}_{\text{el-ph}}|1\rangle|n_{\text{ph}}^{i}\rangle\right|^{2}\times\delta(E_{n_{\text{ph}}}^{f}-E_{n_{\text{ph}}}^{i}-\hbar\omega_{x}),$ (25)
where |$\ket{0}$ and |$\ket{1}$ are the electron’s ground and excited state, |$\ket{n_{p}^{i}}$ and |$\ket{n_{p}^{f}}$ are the phonon’s initial and final states and $|E_{n_{\text{ph}}}^{f}-E_{n_{\text{ph}}}^{i}|=\hbar\omega_{q}$ is the energy of creation of one phonon with the wave vector $\bm{q}$. In addition, although not written explicitly here, this expression needs to be summed up in all possible final phonon states and averaged in all thermally excited initial states.
We write the electron-phonon coupling Hamiltonian via the modulation of dielectric constant as ^$\hat{\mathcal{H}}_{\text{el-ph}}^{(\text{mod})}$ and the displacement of neon surface as ^$\hat{\mathcal{H}}_{\text{el-ph}}^{(\text{dis})}$. The relaxation rate corresponding to each mechanisms and different thermal phonons can be expressed as
$\mathit{\Gamma}^{(\text{dis})}$ $=\frac{2\pi}{\hbar}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\left|\langle n_{q}|\langle 0_{x},0_{y},1_{z}|\hat{\mathcal{H}}^{(\text{dis})}_{\text{el-ph}}|1_{x},0_{y},1_{z}\rangle|n^{\prime}_{q}\rangle\right|^{2}$
×$\times\delta(\sum_{q}\hbar\omega_{q}n_{q}-\sum_{q}\hbar\omega_{q}n^{\prime}_{q}-\hbar\omega_{x}),$
$=\frac{2\pi}{\hbar}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\left|\mathcal{V}^{\text{(dis)}}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)\right|_{10}^{2}G_{\bm{q}_{\scriptscriptstyle\parallel}}$
×$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}),$ (26)
$\mathit{\Gamma}^{(\text{mod})}$ $=\frac{2\pi}{\hbar}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\left|\langle n_{q}|\langle 0_{x},0_{y},1_{z}|\hat{\mathcal{H}}^{(\text{mod})}_{\text{el-ph}}|1_{x},0_{y},1_{z}\rangle|n^{\prime}_{q}\rangle\right|^{2}$
×$\times\delta(\sum_{q}\hbar\omega_{q}n_{q}-\sum_{q}\hbar\omega_{q}n^{\prime}_{q}-\hbar\omega_{x}),$
$=\frac{2\pi}{\hbar}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\left|\mathcal{V}^{\text{(dis)}}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)\right|_{10}^{2}G_{\bm{q}_{\scriptscriptstyle\parallel}}$
×$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}),$ (27)
where $|1_{z}\rangle$ represents the electron’s Rydberg ground state in $z$, $|0_{x},0_{y}\rangle$ and $|1_{x},0_{y}\rangle$ represent the electron’s in-plane ground state and first excited state in $x$. The subscript “10” denotes the excited to ground state transition and $n_{q}=1/[\exp(\hbar\omega_{q}/(k_{\scriptscriptstyle\text{B}}T))-1]\ll 1$ denotes the average thermal occupation number of the phonon state with the wavenumber $q$ at $T=10$ mK. The function $G_{\bm{q}_{\scriptscriptstyle\parallel}}$ is
$G_{\bm{q}_{\scriptscriptstyle\parallel}}$ $=|\langle 0_{x},0_{y},1_{z}|{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}}|1_{x},0_{y},1_{z}\rangle|^{2}$
$=\frac{1}{2}\left(q_{x}a_{x}\right)^{2}{\rm e}^{-(q_{x}^{2}a_{x}^{2}+q_{y}^{2}a_{y}^{2})/2},$ (28)
where $a_{x}=\sqrt{\hbar/m_{e}\omega_{x}}$, $a_{y}=\sqrt{\hbar/m_{e}\omega_{y}}$ are the characteristic trapping widths in $x$ and $y$, respectively.
The phonon-induced dephasing of electrons following the treatment used in ref [32]. The ensemble-averaged mean-square phase increment is given by
[…]
$E_{10}(t)$ $=\langle 1_{x},0_{y},1_{z}|\hat{\mathcal{H}}_{\text{el-ph}}(t)|1_{x},0_{y},1_{z}\rangle$
$-\langle 0_{x},0_{y},1_{z}|\hat{\mathcal{H}}_{\text{el-ph}}(t)|0_{x},0_{y},1_{z}\rangle,$ (31)
is the energy difference between two electronic states. The energy correlator can be approximated by a $\delta$-function which gives the depahsing rate $\Gamma_{\phi}$. Similar to the relaxation calculation, the matrix element for Eq. (31) is given by
$\bra{n_{q}}E_{10}(t^{\prime})E_{10}(t^{\prime\prime})\ket{n_{q}}$ (32)
$=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\left[(\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{11}-(\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{00}\right]^{2}$
×$\times[(n_{q}+1)e^{-i\omega_{q}(t^{\prime}-t^{\prime\prime})}-n_{q}e^{i\omega_{q}(t^{\prime}-t^{\prime\prime})}].$
The integral over time can be found as
[…]
$=\lim_{t\rightarrow\infty}2\pi t\delta(\omega_{q}),$
with $\delta(\omega_{q})=c^{-1}_{\text{Ne}}\delta(q)$ ,which result in
$\mathit{\Gamma}_{\phi}=$ $\frac{2\pi}{\hbar^{2}c_{\text{Ne}}}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\left[(\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{11}-(\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{00}\right]^{2}\delta(q).$ (34)
For one-phonon processes, this summation is zero due to the $\delta$-function, whereas for two-phonon processes, the $\delta$-function becomes $\delta(|q_{1}|-|q_{2}|)$, meaning the annihilation of a phonon with $q_{1}$ and creation one with $q_{2}$. As seen below, since $\mathcal{V}^{\text{(mod)}}$ does not contain $\bm{p}_{\scriptscriptstyle\parallel},\bm{r}_{\scriptscriptstyle\parallel}$, it means $\left|\mathcal{V}^{\text{(mod)}}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)\right|_{11}=\left|\mathcal{V}^{\text{(mod)}}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)\right|_{00}=\text{constant}$. Thus, the sum is essentially determined by $\left[({\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{11}-({\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{00}\right]^{2}=\frac{1}{4}\left(q_{x}a_{x}\right)^{4}{\rm e}^{-(q_{x}^{2}a_{x}^{2}+q_{y}^{2}a_{y}^{2})/2}$. (See supplementary materials of ref [31].) For $\mathcal{V}^{\text{(dis)}}$, only the first term involves $\bm{p}_{\scriptscriptstyle\parallel}$ and contributes to the sum.
We now analyze in detail the two electron-phonon interaction mechanisms.
IV.1 Phonon-induced displacement of neon surface
The electron-phonon coupling strength $\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)$ in Eq. (24) through the mechanism of phonon-induced displacement of neon surface is given by
$\mathcal{V}^{\text{(dis)}}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)$ $=\frac{q_{z}}{q}\sqrt{\frac{\hbar}{V\rho_{\scriptscriptstyle\text{Ne}}c_{\scriptscriptstyle\text{Ne}}q}}\left\{-\frac{{\rm i}}{m_{e}}(\bm{q}_{\scriptscriptstyle\parallel}\cdot\hat{\bm{p}_{\scriptscriptstyle\parallel}})\hat{p}_{z}\right.$
$\left.-\frac{{\rm i}\hbar}{2m_{e}}q_{\scriptscriptstyle\parallel}^{2}\hat{p}_{z}+e\mathcal{E}_{z}+\mathit{\Lambda}q_{\scriptscriptstyle\parallel}^{2}U_{\text{p}}(q_{\scriptscriptstyle\parallel}z)\right\},$ (35)
Here $\hat{\bm{p}_{\scriptscriptstyle\parallel}}=-i\hbar\partial_{\bm{r}}$ is the in-plane electron momentum operator, $\hat{p}_{z}=-i\hbar\partial_{z}$ is the vertical electron momentum operator, $q_{z}/q$ restrict ourselves to the vertical displacement of Ne surface.
The $U_{\text{p}}(\eta)$ function is associated with the change of polarization potential due to surface curvature [33, 34],
$U_{\text{p}}(\eta)=\frac{1}{\eta^{2}}\left[1-\eta K_{1}(\eta)\right].$ (36)
Here $K_{1}(\eta)$ is the modified Bessel function of the first kind. The first two terms describe the kinematic interaction arising from the electron’s wavefunction being set to zero on a non-flat hard-wall surface. The first two terms have nonzero contribution only if we are considering out-of-plane excitation. The pressing electric field term in practical experiments is much smaller than the polarization term. So, we only need to include the polarization term in our calculation.
The relaxation rate can be calculated with Eq. (27),
$\mathit{\Gamma}^{\text{(dis)}}$ $=\frac{2\pi}{\hbar}\frac{\hbar c_{\scriptscriptstyle\text{Ne}}}{V\rho_{\scriptscriptstyle\text{Ne}}}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\frac{q_{z}^{2}}{q^{2}}\frac{1}{qc_{\scriptscriptstyle\text{Ne}}^{2}}G_{\bm{q}_{\scriptscriptstyle\parallel}}U_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$
×$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}),$ (37)
and
$U_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$ $=\left|\mathit{\Lambda}q_{\scriptscriptstyle\parallel}^{2}\langle 1_{z}|U_{\text{p}}(q_{\scriptscriptstyle\parallel}z)|1_{z}\rangle\right|^{2}$
$=\left|\mathit{\Lambda}q_{\scriptscriptstyle\parallel}^{2}\int{\rm d}z\,|\psi_{1}(z)|^{2}\frac{1}{(q_{\scriptscriptstyle\parallel}z)^{2}}\left[1-q_{\scriptscriptstyle\parallel}zK_{1}(q_{\scriptscriptstyle\parallel}z)\right]\right|^{2}$
$\approx-\frac{1}{2}\mathit{\Lambda}^{2}q_{\scriptscriptstyle\parallel}^{4}\ln(q_{\scriptscriptstyle\parallel}r_{\scriptscriptstyle\text{B}}),$ (38)
with the approximation $q_{\scriptscriptstyle\parallel}z\rightarrow 0$ for our problem.
Finally, the decay rate can be written as
$\mathit{\Gamma}^{(\text{dis})}$ $=\frac{R^{2}r_{\scriptscriptstyle\text{B}}^{2}\omega_{0}^{6}}{8\pi m_{e}\rho_{\scriptscriptstyle\text{Ne}}c_{\scriptscriptstyle\text{Ne}}^{9}}\int_{0}^{1}{\rm d}\gamma\,\gamma^{2}(1-\gamma^{2})^{3}$
×$\times\exp\left[-\frac{\hbar\omega_{0}}{2m_{e}c_{\scriptscriptstyle\text{Ne}}^{2}}(1-\gamma^{2})\right]$
×$\times\left[\ln\left(\frac{\omega_{0}}{c_{\scriptscriptstyle\text{Ne}}}r_{\scriptscriptstyle\text{B}}\sqrt{1-\gamma^{2}}\right)\right]^{2}.$ (39)
For simplicity, we have assumed again $\omega_{x}=\omega_{y}=\omega_{0}$. Numerical integral shows $\mathit{\Gamma}^{(\text{dis})}=49.4$ s-1 for the transition frequency at $f_{0}=6.4$ GHz. This leads to the corresponding relaxation time,
$T_{1}^{\left(\text{dis}\right)}=1/\mathit{\Gamma}^{\left(\text{dis}\right)}=20.2\text{ ms},$ (40)
and the coherence time,
$T_{2}^{\left(\text{dis}\right)}=\left(\frac{\mathit{\Gamma}^{\left(\text{dis}\right)}}{2}+\Gamma_{\phi}\right)^{-1}=40.4\text{ ms},$ (41)
both of which are much longer than the reported observations.
Table 1. : Calculated relaxation time $T_{1}^{\left(\text{dis}\right)}$ and coherence time $T_{2}^{\left(\text{dis}\right)}$ of an electron in-plane motional state with varied transition frequency $f_{x}$ for phonon-induced surface displacement.
$f_{x}$ (GHz) $T_{1}^{\left(\text{dis}\right)}$ (s) $T_{2}^{\left(\text{dis}\right)}$ (s)
1 183.1183.1183.1 366.2366.2366.2
2 4.794.794.79 9.589.589.58
3 0.630.630.63 1.261.261.26
4 0.160.160.16 0.320.320.32
5 0.060.060.06 0.120.120.12
6 0.0260.0260.026 $5.31\times 10^{-2}$
7 0.0140.0140.014 $2.81\times 10^{-2}$
8 $8.3\times 10^{-3}$ $1.66\times 10^{-2}$
9 $5.3\times 10^{-3}$ $1.06\times 10^{-2}$
10 $3.6\times 10^{-3}$ $7.29\times 10^{-3}$
Moreover, as seen from Eq. (39), the decay rate $\mathit{\Gamma}^{\left(\text{dis}\right)}$ is dependent on $\omega_{x}$. In Table IV.1 and Fig.2, we present the variations of $T_{1}^{\left(\text{dis}\right)}$ and $T_{2}^{\left(\text{dis}\right)}$ with the transition frequency $f_{x}$. In most cases, $T_{1}^{\left(\text{dis}\right)}$ and $T_{2}^{\left(\text{dis}\right)}$ exceed 1 ms, which surpasses the $T_{1}$ and $T_{2}$ values for most semiconductor as well as superconducting charge qubits.
IV.2 Phonon-induced modulation of dielectric constant
Another decoherence mechanism comes from the phonon-induced modulation of dielectric constant ^$\hat{\mathcal{H}}_{\text{el-ph}}^{(\text{mod})}$. The dielectric constant variation $\delta\epsilon$ and the density variation $\varrho$ is related by
$\delta\epsilon=(\epsilon_{\scriptscriptstyle\text{Ne}}-1)\frac{\varrho}{\rho_{\scriptscriptstyle\text{Ne}}}.$ (42)
The electron-phonon coupling strength $\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)$ in Eq. (24) through the mechanism of phonon-induced modulation of dielectric constant is given by
$\mathcal{V}^{\text{(mod)}}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(z)$ $=-{\rm i}\mathit{\Lambda}q_{\scriptscriptstyle\parallel}\sqrt{\frac{\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}}{V\rho_{\scriptscriptstyle\text{Ne}}c_{\scriptscriptstyle\text{Ne}}^{2}}}\int_{-\infty}^{0}{\rm d}z^{\prime}\,\frac{\sin(q_{z}z^{\prime})}{z-z^{\prime}}$
×$\times K_{1}(q_{\scriptscriptstyle\parallel}|z-z^{\prime}|),$ (43)
which is independent of $\bm{r}_{\scriptscriptstyle\parallel}$ and contains the modified Bessel function of the first kind, $K_{1}(\eta)$.
The relaxation rate can be calculated with Eq. (27),
$\mathit{\Gamma}^{\text{(mod)}}$ $=\frac{2\pi}{\hbar}\frac{\mathit{\Lambda}^{2}}{V\rho_{\scriptscriptstyle\text{Ne}}c_{\scriptscriptstyle\text{Ne}}^{2}}\sum_{\bm{q},q_{z}>0}q^{2}\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}G_{\bm{q}_{\scriptscriptstyle\parallel}}F_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$
×$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}).$ (44)
The function $F_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$ is
$F_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$ $=\left|\langle 1_{z}|\int_{-\infty}^{0}{\rm d}z^{\prime}\,\frac{\sin(q_{z}z^{\prime})}{z-z^{\prime}}K_{1}(q_{\scriptscriptstyle\parallel}|z-z^{\prime}|)|1_{z}\rangle\right|^{2}$
$=\left|\int_{0}^{\infty}{\rm d}z\,|\psi_{1}(z)|^{2}\right.$
×$\left.\times\int_{-\infty}^{0}{\rm d}z^{\prime}\,\frac{\sin(q_{z}z^{\prime})}{z-z^{\prime}}K_{1}(q_{\scriptscriptstyle\parallel}|z-z^{\prime}|)\right|^{2},$ (45)
where $\psi_{1}(z)$ is the normalized electron’s ground-state wavefunction in $z$ given by Eq. (6). To calculate the integrals in $F_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$, since $z>0$ for the electron and $z^{\prime}<0$ for the image charges always hold, $z-z^{\prime}=|z-z^{\prime}|>0$ always holds.
Finally, the decay rate can be written as
$\mathit{\Gamma}^{\text{(mod)}}$ $=\frac{8R^{2}\omega_{0}^{4}}{\pi m_{e}\rho_{\scriptscriptstyle\text{Ne}}c_{\scriptscriptstyle\text{Ne}}^{7}}\int_{0}^{1}{\rm d}\gamma\,(1-\gamma^{2})$
×$\times\exp\left[-\frac{\hbar\omega_{0}}{2m_{e}c_{\scriptscriptstyle\text{Ne}}^{2}}(1-\gamma^{2})\right]$
×$\times\left|\int_{0}^{\infty}{\rm d}s\int_{0}^{\infty}{\rm d}s^{\prime}\,\frac{s^{2}}{(s+s^{\prime})^{2}}\sin\left(\frac{\omega_{0}}{c_{\scriptscriptstyle\text{Ne}}}r_{\scriptscriptstyle\text{B}}us^{\prime}\right){\rm e}^{-2s}\right|^{2},\,$ (46)
where $\gamma\equiv\cos\vartheta$ with $\vartheta$ being the polar angle of phonon wavevectors, $s\equiv z/r_{\scriptscriptstyle\text{B}}>0$, and $s^{\prime}\equiv-z^{\prime}/r_{\scriptscriptstyle\text{B}}>0$. They are all positive dimensionless variables. For simplicity, we have assumed $\omega_{x}=\omega_{y}=\omega_{0}$. Eq. (46) can be numerically integrated without further approximations.
For a qubit transition frequency of $f_{0}=\omega_{0}/2\pi=6.4$ GHz, as commonly used in our experiments, ignoring the practical need of anharmonicity, the numerical integration gives a relaxation rate of $\mathit{\Gamma}^{\text{(mod)}}=908$ s-1. Consequently, the relaxation time $T_{1}$ is
$T_{1}^{\text{(mod)}}=\frac{1}{\mathit{\Gamma}^{\text{(mod)}}}=1.1\,\text{ms}.$ (47)
Hence, the coherence time $T_{2}$ is relaxation-limited by phonon emission [35],
$T_{2}^{\text{(mod)}}=\left(\frac{\mathit{\Gamma}^{\text{(mod)}}}{2}+\mathit{\Gamma}_{\phi}\right)^{-1}=2.2\,\text{ms}.$ (48)
These calculated numbers are at least one order of magnitude better than our experimentally observed $\sim$0.1 ms. Moreover, the decay rate $\mathit{\Gamma}^{(\text{mod})}$ has a complicated dependence on $\omega_{0}$, not only in the prefactor but also in the integrals. In Table 2 and Fig.2 , we present how $T_{1}$ and $T_{2}$ vary with the qubit transition frequency $f_{0}$. In all cases, they are longer than 0.1 ms, and longer than those of most semiconductor charge qubits.
Table 2. : Calculated relaxation time $T_{1}^{\text{(mod)}}$ and coherence time $T_{2}^{\text{(mod)}}$ of an electron in-plane motional state with varied transition frequency $f_{0}$ for phonon-induced dielectric-constant modulation.
$f_{0}$ (GHz) $T_{1}^{\text{(mod)}}$ (s) $T_{2}^{\text{(mod)}}(s)$
1 13.4913.4913.49 27.027.027.0
2 0.330.330.33 0.660.660.66
3 0.0410.0410.041 0.0820.0820.082
4 0.0100.0100.010 $1.98\times 10^{-2}$
5 $3.4\times 10^{-3}$ $6.81\times 10^{-3}$
6 $1.5\times 10^{-3}$ $2.94\times 10^{-3}$
7 $7.4\times 10^{-4}$ $1.48\times 10^{-3}$
8 $4.2\times 10^{-4}$ $8.34\times 10^{-4}$
9 $2.6\times 10^{-4}$ $5.11\times 10^{-4}$
10 $1.7\times 10^{-4}$ $3.33\times 10^{-4}$
So far, we have assumed that the solid Ne occupies the entire lower half space $z<0$. This is a valid approximation when the Ne film is thicker than $\sim$10 nm. However, if the Ne film is only a few nanometers thick, the image charge will be dominated by the underlying dielectric substrate, such as a silicon (Si) or sapphire. Then the electron decoherence is primarily induced by the phonons in Si or sapphire. At low temperatures, the longitudinal sound velocity in silicon and sapphire are $c_{\text{Si}}=8480$ m/s [36] and $c_{\text{Sa}}=11350$ m/s, respectively [37]. Consequently, the typical electron’s in-plane wavenumber $\left(m\omega_{x}/\hbar\right)^{1/2}=1.9\times 10^{7}$ m-1 is much larger than the wavenumbers of the resonant phonons $\omega_{0}/c_{\text{Si}}=4.7\times 10^{6}$ m-1 and $\omega_{0}/c_{\text{Sa}}=3.5\times 10^{6}$ m-1 ($f_{0}=6.4$ GHz). This mismatch makes the relaxation rate through these channels essentially 0, following the above calculation. Physically it originates from an inability to simultaneously satisfy energy and momentum conservation during the scattering process. Therefore, the single phonon scattering from the substrate is exponentially suppressed due to the mismatch between the size of the electron wave function and the phonon wavelength at the same energy. In experiments, the thickness of the Ne film typically lies between the two aforementioned extremes. As a result, the actual electron charge coherence time may exceed our conservative estimations above.
Fig. 2. : Calculated relaxation time $T_{1}$ (top) and coherence time $T_{2}$ (bottom) for the two electron-phonon coupling mechanisms: phonon-induced surface displacement and phonon-induced dielectric-constant modulation.
V Conclusion
In our comprehensive theoretical study, we examined the decoherence mechanisms impacting the motional states of a single electron on a solid Ne surface at a temperature of 101010 mK. Notably, we found that the relaxation time for electron motional states with a transition frequency of 6.4 GHz is $T_{1}^{\left(\text{vac}\right)}=\,99$ s in vacuum and $T_{1}^{\left(\text{cav}\right)}=\,32$ ms in a superconducting cavity with coupling constant $g\sim 2\pi\times 5$ MHz, resonator linewidth $\kappa\sim 2\pi\times 0.5$ MHz, and detuning from resonant frequency $\left|f_{x}-f_{\text{r}}\right|\sim 500$ MHz. The coherence time $T_{2}^{\left(\text{vac}\right)}$ has upper bounds of 198 s and 64 ms, respectively. Additionally, we investigated the decoherence processes caused by phonon-induced displacement of the Ne surface, finding coherence times $T_{2}^{\left(\text{dis}\right)}$ ranging from approximately 366 s to 7 ms, spanning the transition frequency from 1 GHz to 10 GHz. Furthermore, we discovered that phonon-induced modulations of Ne relative permittivity can significantly alter the coherence time $T_{2}^{\left(\text{mod}\right)}$ from 27 s to 0.3 ms, within the same transiton frequency range. These results highlight the promising potential of single-electron motional (charge) qubits on solid Ne that may outperform most existing semiconductor and superconducting charge qubits.
Acknowledgements.
D. J., X. L., and Q. C. acknowledge support from the Argonne National Laboratory Directed Research and Development (LDRD) program. D. J., X. L., and S. Z. acknowledge support from the Air Force Office of Scientific Research (AFOSR) under Award No. FA9550-23-1-0636. D. J. acknowledges support from the Department of Energy (DOE) under Award No. DE-SC0025542 and the National Science Foundation (NSF) under Award No. OSI-2426768. Work performed at the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, was supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DEAC02-06CH11357. The authors thank Mark I. Dykman for helpful discussions.